OpenStudy (anonymous):

The points (3, 24) and (7, 56) represent points of a function where y, the number of photographs, varies directly with x, the number of pages in an album. Which statement describes another point on the graph of this function?

OpenStudy (anonymous):

We could find the slope and create an equation. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{56-24}{7-3}\) Subtract those.

OpenStudy (anonymous):

Can you do that? @Olayshia22

OpenStudy (anonymous):

Yes i got it thanks

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\) You can give medals by clicking 'Best Resposne'.

OpenStudy (anonymous):

*Response